OpenStudy (anonymous):

What is the pH of a 6.40 × 10-3 M solution of sodium hydroxide, NaOH?

OpenStudy (anonymous):

What is the pH of a 6.40 × 10-3 M solution of sodium hydroxide, NaOH?

OpenStudy (anonymous):

If 20.0 mL of a 0.500 M HCl solution is required to neutralize 35.0 mL KOH, what is the concentration of the KOH solution?

OpenStudy (anonymous):

For the first problem, you should take [OH-] = 6.40 x 10^(-3) since NaOH will dissociate completely. Then you can find pOH = -log[6.40 x 10^(-3)] = 2.1938 After you have pOH, plug in to: pOH + pH = 14, and solve for pH. 2.1938 + pH = 14 pH = 11.8, a basic solution. This makes sense because you have a fairly concentrated base in NaOH. ======================================== For the second problem, you need to use the formula \[M _{1}V _{1} = M _{2}V _{2}\], where 1 is represented by HCl and 2 is represented by KOH: (0.500 M)(20.0 ml) = x(35.0 ml) x = 0.286 M, a lower concentration than the HCl, which makes sense because you have more of the KOH to begin with.

OpenStudy (anonymous):

Consider the following reaction: 4NH3 (g) + 5O2 (g) 4NO (g) + 6H2O (g)